# Pre-Chirp-Domain Index Modulation for Full-Diversity Affine Frequency Division Multiplexing towards 6G

### v3

@@ VII Conclusions
The corresponding two columns in the matrix $\mathbf{\Phi}(\delta)$ can be expressed as (64) and (65) at the top of the next page, where $\hat{H}_{(\cdot)}[\cdot]\hat{x}_{(\cdot)}$ and $H_{(\cdot)}[\cdot]x_{(\cdot)}$ represent the corresponding elements in $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ and $\mathbf{\Phi}(\mathbf{x})$, respectively.
@@
$\hskip 28.45274pt=\sum_{p=1}^{P}\sum_{n=1}^{N}\left|e^{\textsf{j}\frac{2\pi}{N}\left[Nc_{2,\operatorname{loc}_{p}+n}^{\prime}\left(\operatorname{loc}_{p}+n\right)^{2}-{N}c_{2,n}^{\prime}n^{2}\right]}x^{\prime}_{\operatorname{loc}_{p}+n}]-e^{\textsf{j}\frac{2\pi}{{N}}\left[{N}c_{2,\operatorname{loc}_{p}+n}\left(\operatorname{loc}_{p}+n\right)^{2}-{N}c_{2,n}n^{2}\right]}x_{\operatorname{loc}_{p}+n}]\right|^{2}\!\!.\!$ (74)
@@
Therefore, the norm of $\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})-\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}$ in (47) can be calculated as (-A), shown at the bottom of this page.
@@
$\hskip 8.53581pt=\!\sum_{p=1}^{P}\sum_{n=1}^{N}\!2\!\left(1\!-\!\mathfrak{R}\left(x_{\operatorname{loc}_{p}+n}^{\prime}\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}e^{\textsf{j}(\theta_{n}^{\prime}-\theta_{n})}\right)\!\right)\!,\!$ (76)
where $x_{\mathrm{loc}_{p}+n}^{\prime}$ and $x_{\mathrm{loc}_{p}+n}$ are the elements of $\mathbf{x}^{\prime}$ and $\mathbf{x}$, respectively, and $\theta_{n}^{\prime}$ and $\theta_{n}$ are given in (51).
@@
$\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})\!-\!\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}\!=\!\sum_{p=1}^{P}\!\sum_{n=1}^{N}\!1\!-\!\cos\left(\psi_{n}\!+\!\theta_{n}^{\prime}\!-\!\theta_{n}\right),$ (77)
where $\psi_{n}$ is defined in (50), i.e., it represents the phase difference between $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$. Substituting (77) into (47) leads to $O_{k,j}(\mathcal{P}_{c})$ of (49).

### v4

@@ VII Conclusions
The corresponding two columns in the matrix $\mathbf{\Phi}(\delta)$ can be expressed as (70) and (71) at the top of this page, where $\hat{H}_{(\cdot)}[\cdot]\hat{x}_{(\cdot)}$ and $H_{(\cdot)}[\cdot]x_{(\cdot)}$ represent the corresponding elements in $\hat{\mathbf{\Phi}}(\hat{\mathbf{x}})$ and $\mathbf{\Phi}(\mathbf{x})$, respectively.
@@
$\hskip 28.45274pt=\sum_{p=1}^{P}\sum_{n=1}^{N}\left|e^{\textsf{j}\frac{2\pi}{N}\left[Nc_{2,\operatorname{loc}_{p}+n}^{\prime}\left(\operatorname{loc}_{p}+n\right)^{2}-{N}c_{2,n}^{\prime}n^{2}\right]}x^{\prime}_{\operatorname{loc}_{p}+n}]-e^{\textsf{j}\frac{2\pi}{{N}}\left[{N}c_{2,\operatorname{loc}_{p}+n}\left(\operatorname{loc}_{p}+n\right)^{2}-{N}c_{2,n}n^{2}\right]}x_{\operatorname{loc}_{p}+n}]\right|^{2}\!\!,\!$ (80)
@@
Therefore, the norm of $\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})-\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}$ in (50) can be calculated as (-A), shown at the bottom of last page, where $x_{\mathrm{loc}_{p}+n}^{\prime}$ and $x_{\mathrm{loc}_{p}+n}$ are the elements of $\mathbf{x}^{\prime}$ and $\mathbf{x}$, respectively. We define $x_{\mathrm{loc}_{p}+n}=r_{p,n}e^{j\vartheta_{p,n}}$ and $x_{\mathrm{loc}_{p}+n}^{\prime}=r_{p,n}^{\prime}e^{j\vartheta_{p,n}^{\prime}}$.
@@
$\hskip 8.53581pt=\!\sum_{p=1}^{P}\sum_{n=1}^{N}\!\!\left(\left|x_{\mathrm{loc}_{p}+n}\right|^{2}+\left|x_{\mathrm{loc}_{p}+n}^{\prime}\right|^{2}\!\right.$
$\hskip 56.9055pt\left.-2\mathfrak{R}\left(x_{\operatorname{loc}_{p}+n}^{\prime}\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}e^{\textsf{j}(\theta_{n}^{\prime}-\theta_{n})}\right)\!\right)\!$
[…]
$\hskip 56.9055pt\left.-2r_{p,n}r_{p,n}^{\prime}\mathfrak{R}\left(e^{\textsf{j}\left(\theta_{n}^{\prime}-\theta_{n}+\vartheta_{p,n}^{\prime}-\vartheta_{p,n}\right)}\right)\!\right)\!,\!$ (82)
where $\theta_{n}^{\prime}$ and $\theta_{n}$ are given by (52).
@@
$\left\|(\mathbf{\Phi}_{k}^{r}(\mathbf{x}^{\prime})\!-\!\mathbf{\Phi}_{j}^{r}(\mathbf{x}))\right\|_{\mathrm{F}}^{2}\!\hskip 8.53581pt=\sum_{p=1}^{P}\sum_{n=1}^{N}F_{p,n},$ (83)
where $F_{p,n}$ is given by (53). Substituting (83) into (50) leads to $O_{k,j}(\mathcal{P}_{c})$ of (54).
-C Relationship between Speed and Normalized Doppler Shift
Let the maximum speed of the mobile station be $v_{e}$, the carrier frequency be $f_{c}$, and the speed of light be denoted by $c$. Then the maximum Doppler frequency $\nu_{\max}$ is given by
$\nu_{\max}\;=\;\frac{v_{e}\,f_{c}}{c}.$ (84)
As established in Section II, the maximum normalized Doppler shift is obtained through the process of normalization as
$\alpha_{\max}\!$ $=\!N\,T\,\nu_{\max}=\frac{\nu_{\max}}{f_{s}}=\frac{v_{e}\,f_{c}}{c\,f_{s}}.$ (85)
where $f_{s}$ represents the subcarrier spacing.